A photographer has two bottles of developer, one a 70% solution, the other, 20% solution. How much developer should he take from each bottle to make 10 ounces of 50% solution?
Answer by richwmiller(17219)

a+b=10,
0.7*a+0.2*b=0.5*10
a=10-b
0.7*(10-b)+0.2*b=5
7-0.7b+0.2*b=5
-0.5*b=-2
b=4
a=10-b
a=6 oz. at 70%
b=4 oz. at 20%
check
0.7*6+0.2*4=0.5*10
4.2+0.8=5
5=5
ok
codemx1